OBSERVATIONS OF DIFFUSE INTERSTELLAR BANDS NEAR 680 nm

\underline{H. TADA}, K. KAWAGUCHI, Department of Chemistry, Faculty of Science, Okayama University, Okayama 700-8530, Japan; H. IZUMIURA, Okayama Astrophysical Observatory, Asaguchi, 719-0232, Japan; T.. SEDIVCOVA and S. CIVIS, J.Heyrovsky Institute of Physical Chmistry, 18223 Prague, Czech Republic.

There is a report on a series of diffuse interstellar bands (DIBs) with a regular spacing of about 35 cm^{−1} near 680 nm^{a}. Glinski et al. proposed that the band is due to a perpendicular vibronic band of a CH_{2}X type molecule, where X is a heavy atom or a linear carbon chain^{b} ^{c} . The K-sub bands appear every 2A, where A is the rotational constant. If the 2A is 17-18 cm^{−1}, the 35 cm^{−1} interval is explained by transitions among the ortho states. To examine the CH_{2}X hypothesis, we observed this series of DIBs toward HD 21389, HD 41117 and HD 198478 with a resolution of 75000 at the Okayama Astrophysical Observatory. We detected the series of DIBs with 35 cm^{−1} intervals. However, spectra of para states are not detected clearly, although the intensity is expected to be one third of that of the ortho states. The observational result is discussed with the simulated spectra of the CH_{2}X type molecules.

aG. H. Herbig, Astrophys. J. \underline{331}, 999 1988

bR. J. Glinski, and J. A. Nuth Publ. Astron. Soc. Pacific \underline{107}, 453 1995

cS. A. Schulz, J. E. King, and R. J. Glinski Mon. Not. R. Astron. Soc. \underline{312}, 769 2000